find "f'(3)" for"f(x) = x\u00b2 + 10x"

1
Expert's answer
2021-01-28T07:11:53-0500

"f(x)=x^2+10x"

"\implies f'(x)=\frac{d}{dx}(f(x))=\frac{d}{dx}(x^2+10x)=2x+10"

And "f'(3)=2(3)+10=6+10=16"